OpenStudy (anonymous):

What are the coordinates of the vertex of the graph of the function y = –3x² –12x + 3?

OpenStudy (anonymous):

huuurry C;

OpenStudy (campbell_st):

there are a couple of ways to do this 1. find the line of symmetry, the vertex lies on the line for a quadratic \[ax^2 +bx + c\] the line of symmetry is \[x = \frac{-b}{2a} \] so in your question \[x = \frac{12}{2 \times -3} ... so.. x = -2\] substitute it into the equation for find the y value.. 2. take out the common factor \[y = -3(x^2 + 4x - 1)\] complete the square on \[x^2 + 4x \] will give \[y = -3((x^2 + 4x + 4) -5)\] so \[y = -3(x + 2) + 15 \] the general form is y = (x -h)^2 + k so the vertex is (h, k) both methods give the same answer.

OpenStudy (anonymous):

Thaank You So Muuch C;